OpenStudy (bbb911):

this is Geometry: Suppose <1 and <2 are vertical angles. What is the value of y if m <1 = y + 82 and m <2 = 6y-8?

OpenStudy (anonymous):

I don't understand your notation. I assume that <1 and <2 are angles, but I'm not familiar with the term 'vertical angle' or with the equation m <1 = y + 82

OpenStudy (anonymous):

etc

OpenStudy (bbb911):

so you dont know it? :(

OpenStudy (anonymous):

I probably do. I don't understand what you're asking.

OpenStudy (bbb911):

oh all of these: < are angles

OpenStudy (anonymous):

I get that. What do you mean by m<1 ?

OpenStudy (bbb911):

well i just copied whatever the question had and it said what is the value of y if m <1 =y+82 and m < 2 =6y-8

OpenStudy (anonymous):

When you say 'vertical angle' do you mean perpendicular?

OpenStudy (bbb911):

vertical angles are the angles opposite each other when two lines cross

OpenStudy (anonymous):

m<1 means measure of <1. What is the magnitude of the angle

OpenStudy (anonymous):

Ah. Thank you.

OpenStudy (anonymous):

vertical angles are congruent so there measures are the same so set the equations equal and solve

OpenStudy (anonymous):

their*

OpenStudy (bbb911):

so are you still going to help me